OpenStudy (anonymous):

what is the value of sin45 degree and cos 45 degree?? i think sin45= 1/underroot 2 and cos45=1/underroot 2 but open the link below here is what value??

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

there are three ways to write the answer one is \[\sqrt{\frac{1}{2}}\] the square root of one half you never see that

OpenStudy (anonymous):

the other is \[\frac{1}{\sqrt2}\] which you sometimes see same thing

OpenStudy (anonymous):

mostly you see it with the denominator rationalized as \[\frac{\sqrt2}{2}\]

OpenStudy (anonymous):

yes but is it wrong?

OpenStudy (anonymous):

they are all the same number

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

how is it the same number?

OpenStudy (anonymous):

@mathmath333 and @satellite73 if both has same values then can i use always \(\dfrac{1}{sqrt2}\)

OpenStudy (mathmath333):

yes